OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

there is no x value as there is no solution to the system; solve 3x + y = 6 for y, getting y = 6 - 3x; subbing into the second equation we see 6x + 2(6-3x) = 8 is impossible, since when we simplify 6x + 12 - 6x = 8, and further: 12 = 8, which is a contradiction.